There are 4 consecutive multiples of 3 such that the sum of the first 3 multiples is twice the fourth. Find the four multiples of 3
Answer by Edwin McCravy(20081)

x = 1st multiple of 3
x+3 = 2nd multiple of 3
x+6 = 3rd multiple of 3
x+9 = 4th multiple of 3

1st + 2nd + 3rd = 2*4th

 x  + x+3 + x+6 = 2(x+9)
           3x+9 = 2x+18
              x = 9

x   = 1st multiple of 3 = 9
x+3 = 2nd multiple of 3 =  9+3 = 12
x+6 = 3rd multiple of 3 = 12+3 = 15
x+9 = 4th multiple of 3 = 15+3 = 18
 

Edwin